OpenStudy (liliakarina):

How to transform rectangular points to polar? E. g., A(-3, -3)? And how to do it from polar to rectangular? E. g., B(2, 3/2pi)?

OpenStudy (anonymous):

to change rectangular to polar in A(-3, -3), use the fact that : \[\large r=\sqrt{x^2+y^2} \] \[\large \theta=arctan(\frac{y}{x}) \] to change from polar to rectangular in \(\large (2,\frac{3\pi}{2}) \), use: \[\large x=rcos\theta \] \[\large y=rsin\theta \]

OpenStudy (anonymous):

in A(-3, -3): r = sqrt(18) = 3sqrt2 theta = arctan(-3/-3) = 5pi/4 so in polar form it's (3sqrt2, 5pi/4)

OpenStudy (liliakarina):

Thank you. I get it now.